# How to Bridge the Sim-to-Real Gap in Digital Twin-Aided Telecommunication Networks

### v1

@@ IV-C Context-Aware Prediction-Powered Inference
Context-aware PPI tackles the lack of contextuality of conventional PPI by allowing the bias correction applied to the training loss to vary with context. The core idea is to partition the synthetic data, and corresponding real validation data, as a function of relevant context features, and then perform bias estimation separately for each context. For instance, one can assign different correction factors for different ranges of a radio channel condition metric or for different cell sectors in a network. By doing so, the model treats synthetic data from contexts where the twin is reliable with more trust, and discounts or heavily corrects synthetic data in contexts where the twin is less accurate [11].
In Fig. 6, following [11], we consider an outdoor urban environment with a single base station (BS), where the propagation paths between the BS and the devices are generated using ray‑tracing as in [12]. Each device position is associated with a ground‑truth azimuth and elevation angle of departure (AoD), which is determined by selecting the propagation path with the highest received power. The objective is to train a mapping that outputs an estimate of the optimal AoD at each receive location, given a real dataset and a much larger synthetic dataset. Each data point is augmented with a binary context variable indicating if the device is within the LoS of the BS.
Panels (a) and (b) present geographic heatmaps of beamforming directions: panel (a) shows the ground truth obtained from real data, while panel (b) displays the prediction made by a DT. Notably, substantial discrepancies emerge in the non-line-of-sight (NLoS) region, where the DT fails to capture the true beamforming pattern, yielding a large sim-to-real gap. Panels (c)–(e) quantify the beam-pointing error for three training schemes: ERM, which relies solely on real data, PPI [13, 4], and context-aware PPI (CPPI) [11], respectively. As seen, PPI performs poorly in the NLoS locations, exhibiting even worse performance than ERM. In contrast, CPPI leverages context-dependent weightings of synthetic samples, thus achieving uniformly lower errors across the entire region.
@@ V Conclusion
The techniques discussed in this paper are key to designing AI models that harness the full power of modern machine learning-based simulation tools, while maintaining reliability in the face of real-world network complexities. However, several challenges remain. Network non-stationarity demands adaptive calibration, where approaches like meta-learning frameworks [14] may be highly valuable. Furthermore, distributional shifts can erode uncertainty estimates, making the integration of techniques such as adaptive sampling [15] crucial to maintaining the robustness of PPI. Finally, generating synthetic samples within the DT, whether via ray tracing or ML‑based emulation, incurs tangible costs. Carefully balancing the volume of synthetic data against the budgetary constraints and the targeted model performance is essential for successful deployment.

### v2

@@ IV-C Context-Aware Prediction-Powered Inference
Context-aware PPI tackles the lack of contextuality of conventional PPI by allowing the bias correction applied to the training loss to vary with context. The core idea is to partition the synthetic data, and corresponding real validation data, as a function of relevant context features, and then perform bias estimation separately for each context. For instance, one can assign different correction factors for different ranges of a radio channel condition metric or for different cell sectors in a network. By doing so, the model treats synthetic data from contexts where the twin is reliable with more trust, and discounts or heavily corrects synthetic data in contexts where the twin is less accurate [ruah2025context].
In Fig. 6, following [ruah2025context], we consider an outdoor urban environment with a single BS, where the propagation paths between the BS and the devices are generated using ray‑tracing as in [zeng2021toward]. Each device position is associated with a ground‑truth azimuth and elevation angle of departure (AoD), which is determined by selecting the propagation path with the highest received power. The objective is to train a mapping that outputs an estimate of the optimal AoD at each receive location, given a real dataset and a much larger synthetic dataset. Each data point is augmented with a binary context variable indicating if the device is within the LoS of the BS.
Panels (a) and (b) present geographic heatmaps of beamforming directions: panel (a) shows the ground truth obtained from real data, while panel (b) displays the prediction made by a DT. Notably, substantial discrepancies emerge in the non-line-of-sight (NLoS) region, where the DT fails to capture the true beamforming pattern, yielding a large sim-to-real gap. Panels (c)–(e) quantify the beam-pointing error for three training schemes: ERM, which relies solely on real data, PPI [angelopoulos2023prediction], and context-aware PPI (CPPI) [ruah2025context], respectively. As seen, PPI performs poorly in the NLoS locations, exhibiting even worse performance than ERM. In contrast, CPPI leverages context-dependent weightings of synthetic samples, thus achieving uniformly lower errors across the entire region.
@@ V Conclusion
We conclude by briefly discussing challenges and further topics.
Real-time DTs. When DTs are used for real-time tasks, a fundamental tension exists between DT fidelity and computational efficiency: high-fidelity simulations provide accurate data but may be too slow for time-critical applications, while fast emulation models sacrifice accuracy [hashash2022edge, pegurri2024toward].
Cross-site transferability and domain adaptation. Calibrated DTs are inherently site-specific, and the application of a calibrated DT from one location or configuration to another typically requires the use of transfer learning and domain adaptation techniques.
Continuous adaptation. In practice, the design and calibration of a DT for data augmentation must be updated either periodically or in response to changes in the environment, benefiting from frameworks such as continual learning and meta-learning [simeone2022machine, chen2023learning].
Clement Ruah (Member, IEEE) is a Research Associate at King’s Laboratory for Intelligent Computing. He received his Ph.D. degree in Information Engineering from King’s College London in 2025. His research focuses on reliable machine learning and mixing simulation-based approaches with uncertainty-aware data-driven solutions.
Houssem Sifaou (Member, IEEE) received the M.S. and Ph.D. degrees in Electrical Engineering from the King Abdullah University of Science and Technology (KAUST) in 2016 and 2021, respectively. From 2021 to 2023, he served as a Research Associate at Imperial College London. He is currently a Research Associate at King’s College London. His research lies at the intersection of machine learning and wireless communications.
Osvaldo Simeone (Fellow, IEEE) is a Professor of Information Engineering at King’s College London, where he co-directs the Centre for Intelligent Information Processing Systems and leads the King’s Communications, Learning and Information Processing lab. He is also a Visiting Professor at Aalborg University. He has received multiple best paper and tutorial awards, and his research has been supported by the NSF, ERC, EPSRC, European Commission, ESA, and industry. Prof. Simeone is a Fellow of the IEEE, IET, and EPSRC, and the author of two textbooks published by Cambridge University Press.
Bashir M. Al-Hashimi worked in the electronics design industry for eight years prior to embarking on an academic career with the School of Electronics and Computer Science, University of Southampton, in 1999. He is currently Vice President (Research & Innovation) at King’s College London and co-Director of the King’s Institute for Artificial Intelligence. He has successfully led a number of large-scale interdisciplinary research programmers funded by the EPSRC and industry, supervised 40 Ph.D. students to successful completion, published 380 refereed technical papers, and authored or coauthored seven books.